# CLAP: Unsupervised 3D Representation Learning for Fusion 3D Perception via Curvature Sampling and Prototype Learning

### v1

@@ Abstract
Unsupervised 3D representation learning via masked-and-reconstruction with differentiable rendering is promising to reduce the labeling burden for fusion 3D perception. However, previous literature conduct pre-training for different modalities separately because of the hight GPU memory consumption. Consequently, the interaction between the two modalities (images and point clouds) is neglected during pre-training. In this paper, we explore joint unsupervised pre-training for fusion 3D perception via differentiable rendering and propose CLAP, short for Curvature sampLing and swApping Prototype assignment prediction. The contributions are three-fold. 1) To overcome the GPU memory consumption problem, we propose Curvature Sampling to sample the more informative points/pixels for pre-training. 2) We propose to use learnable prototypes to represent parts of the scenes in a common feature space and bring the idea of swapping prototype assignment prediction to learn the interaction between the two modalities. 3) To further optimize learnable prototypes, we propose an Expectation-Maximization training scheme to maximize the similarity between embeddings and prototypes, followed by a Gram Matrix Regularization Loss to avoid collapse. Experiment results on NuScenes show that CLAP achieves $300\%$ more performance gain as compared to previous SOTA 3D pre-training method via differentiable rendering. Codes and models will be released.
@@ 1 Introduction
3D perception provides understanding of the ego vehicle’s surrounding 3D space and plays an important role in downstream prediction and control tasks in autonomous driving. Single-modality 3D perception has long been studied with camera [11, 6, 40, 45, 55, 19] or LiDAR (Light-Detection-And-Ranging) [36, 50, 39, 54, 12, 1] inputs, which is less robust to changing lighting and weather conditions [30]. To compensate shortages of different modalities, research prevails on fusion 3D perception with camera images and LiDAR point clouds [27, 25, 21, 24, 20, 31, 7].
@@
Figure 1: Unlike previous unsupervised 3D representation learning method UniPAD [53] that separately pre-train LiDAR and camera encoders (a), our proposed method CLAP conducts joint pre-training for fusion perception.
Meanwhile, labeling in 3D space is notoriously time-and-energy-consuming. Unsupervised 3D representation learning [23, 18, 5, 52, 48, 17, 57], which pre-trains backbones without any label and applies the pre-trained weight to initialize downstream models for performance improvement, shows potential to alleviate labeling burden in 3D perception. Hence, the community starts to explore pre-training for fusion 3D perception in an unsupervised manner [41, 56, 22, 53], among which UniPAD [53] shows superior performance. UniPAD [53] is a masked-reconstruction-based method and utilizes differentiable rendering as decoder during pre-training. However, during pre-training stage, the camera and LiDAR encoders are separately pre-trained without considering the interaction between the two modalities in [53]. The reason stems from the high GPU memory consumption. If all of the LiDAR points and camera pixels are considered for pre-training, the most advanced GPU up-to-date (H100) is only able to hold batchsize=1. But batchsize is an important hyperparameter for image encoder and a small one normally brings degradation to the performance, making it unfeasible for 3D pre-training.
In this paper, we explore joint unsupervised pre-training for fusion 3D perception via differentiable rendering and propose CLAP, which is shorthand for Curvature sampLing and swApping Prototype assignment prediction. To begin with, we need to sample the LiDAR points and camera pixels for pre-training in order to make joint pre-training feasible. The most direct way is to use uniform sampling but as the sample number is small compared to the raw inputs ($\sim\frac{1}{100}$), uniform sampling does not brings improvement against separate pre-training. In order to sample more informative parts of the environment for pre-training, CLAP first estimates curvature of each LiDAR point in the 3D space by taking second order derivative of the SDF (signed distance field) function. Then we compute weights proportional to the estimated curvature and sample more informative points according to these weights. As for pixel sampling, we project the LiDAR points onto image planes and assign the weights to pixel. As LiDAR point clouds are sparse, we further apply gaussian blur to broadcast the weights to nearby pixels and sample pixels according to these weights. With Curvature Sampling, we further consider how to better utilize the interaction between camera images and LiDAR point clouds to gain some level of understanding about “objectness” or “parts of objects” in an unsupervised way. Inspired by SwAV [3], we propose to use learnable prototypes to learn to represent parts of the scenes. Similar to [3], CLAP applies a swapping prototype assignment prediction loss between camera and LiDAR embeddings to learn the common feature space. To further optimize the learnable prototypes, an Expectation-Maximization training scheme is proposed. In the Expectation step, we compute the probability that each prototype is assigned to the 3D embeddings in the 3D space. Then in the Maximization step, we maximize the similarity between embeddings and prototypes by minimizing the entropy of the assignment scores. Finally, to avoid collapse of the prototype [3], a gram matrix regularization Loss is proposed to minimize the correspondence among prototypes.
Through extensive experiments on the popular autonomous driving dataset NuScenes [2], we demonstrate that CLAP is able to achieve up to $300\%$ more improvement than previous SOTA method [53].
@@ 2 Related Work
Fusion 3D Object Detection. Light-Detection-And-Ranging (LiDAR) and camera are important sensors for autonomous driving perception. Previous works mainly focus on single-modality 3D perception. For LiDAR-based 3D object detection, there are three main streams with different embedding schemes for point clouds inputs. 1) Embraced by [36, 38], point-based methods utilize point-level embeddings for 3D object detection. 2) Voxel-based methods [50, 10, 54, 1, 51, 12] voxelize the 3D scene and use sparse convolution or transformer for embedding. 3) Point-voxel-combined methods [37, 39] utilize both embeddings from 1) and 2). For camera-based 3D perception, [11, 6, 40, 45, 55, 19] embed 2D features on image plane and project these 2D features into 3D space with estimated depth. As single-modality perception can be degraded by various weather and lighting conditions [30], fusion 3D perception [7, 27, 25, 21, 24, 20, 31] starts to flourish. These methods mainly focus on the way to integrate embeddings of different modalities and use supervised pre-trained 2D and 3D backbones to achieve the SOTA performance. As labeling in 3D space is costly, we explore unsupervised 3D representation learning for fusion 3D perception.
@@
3D Pre-training. Annotating for 3D data is notoriously time-and-energy-consuming and the emergence of unsupervised representation learning for 2D image [14, 43, 3, 13, 46, 15] provides a promising way to alleviate the annotation burden. Researcher starts to introduce unsupervised 3D representation learning into scene-level 3D point clouds [47, 26, 16, 5, 18, 23, 34, 52, 48, 17, 57], which can be divided into two contrastive-based and masked-and-reconstruction-based branches. Embraced by [47, 26, 16, 5, 18, 23, 34], contrastive-based works propose various ways to build suitable views and conduct contrastive learning to improve the performance in downstream perception task. Inspired by [15] in image domain, [52, 48, 17, 57] propose to first mask the input point clouds and pre-train the 3D encoders with a shallow decoder for reconstructing the unmasked inputs. [41, 56, 22] are pioneering works to introduce contrastive learning into fusion perception. They consider camera and LiDAR embeddings as different views of the scene and apply the contrastive loss between the two modalities. Inspired by the success of neural field in representing 3D scenes [32, 44] and previous attempts to introduce neural rendering to 3D pre-training for point clouds [49, 17, 57], UniPAD [53] proposes to use a differentiable-rendering decoder for masked-and-reconstruction pre-training and achieves SOTA performance for unsupervised 3D representation learning on fusion 3D perception. However, due to the high GPU memory consumption, UniPAD [53] is only able to separately pre-train the image and point cloud encoders and fails to utilize the interaction between modalities during pre-training. In this paper, we explore unsupervised joint pre-training for 2D and 3D backbones via differentiable rendering.
@@ 3 Method
Figure 2: The pipeline of CLAP. In order to jointly pre-train the LiDAR, camera and fusion encoders, we first embed the paired LiDAR point clouds and camera images with $f^{\text{enc}}_{\text{P}}$, $f^{\text{enc}}_{\text{I}}$ and $f^{\text{enc}}_{\text{fusion}}$. Then based on the fusion features, CLAP applies differentiable rendering to predict both depth and rgb with the SDF and RGB values of the sampled points along LiDAR/camera rays from $f^{\text{SDF}}$ and $f^{\text{RGB}}$, with which we compute loss against the observed LiDAR point cloud and camera images. To make joint pre-training feasible, we propose Curvature Sampling to sample informative parts of the 3D scene, as descriped in (a). Furthermore, we propose to use learnable prototypes to represent parts of objects and use a swapping prototype assignment prediction loss to incorporate the interaction of different modalities into pre-training. Finally, we optimize the learnable prototypes with an Expectation-Maximization scheme together with a gram matrix regularization loss.
In this section, we introduce CLAP for unsupervised 3D pre-training via differentiable rendering on fusion 3D perception. As described in Fig. 2, CLAP pre-trains the image, LiDAR and fusion encoders jointly with Neural Field Rendering. In order to make joint pre-training feasible, Curvature Sampling is proposed as shown in Fig. 2 (a). To further make use of both modalities, we propose a swapping prototype assignment prediction scheme, which is visualized in Fig. 2 (b). We first discuss the formulation and overall pipeline in Section 3.1. Then we introduce the details about neural field and differentiable rendering in Section 3.2. Finally, we describe the curvature sampling and swapping prototype assignment prediction separately in Section 3.3 and 3.4.
@@ 3.1 Formulation and Pipeline
Notations. To begin with, we denote the input image set from $N_{\text{cam}}$ cameras as =$\mathcal{I}=\{\mathbf{I}_{n}\in\mathbb{R}^{H\times W\times 3}\}_{n=1}^{N_{\text{cam}}}$ and LiDAR point cloud as $\mathbf{P}\in\mathbb{R}^{N_{p}\times(3+d)}$. $H$ and $W$ are the height and width of the images and each pixel on the images has $3$ values for RGB. $N_{p}$ is the number of points in the LiDAR point cloud and each of them contains $xyz$-location and $d$ feature channels. For example, in NuScenes [2] dataset, $d=2$ represents the intensity and timestamp of each point and there are $N_{\text{cam}}=6$ surrounding cameras on the autonomous vehicle. For each pair of camera image and LiDAR point cloud, we have the transformation matrix $\mathbf{T}_{n}\in\mathbb{R}^{3\times 4}$ indicating the projection between each camera plane and LiDAR coordinate, where $n=1,2,...,N_{\text{cam}}$.
@@
where $\hat{\mathbf{P}}\in\mathbb{R}^{\hat{D}\times\hat{H}\times\hat{W}\times\hat{d}_{\text{P}}}$ is the embedded 3D features for LiDAR point cloud. $\hat{D}$, $\hat{H}$ and $\hat{W}$ are spatial resolutions of the embedded features and $\hat{d}_{\text{P}}$ is number of feature channels after encoding. Then for camera images $\mathcal{I}$, we encode them and project the 2D features to 3D space with $\mathcal{T}$ with image encoder $f^{\text{enc}}_{\text{I}}$
$\hat{\mathbf{I}}=f^{\text{enc}}_{\text{I}}(\mathcal{I},\{\mathbf{T}_{n}\}_{n=1}^{N_{\text{cam}}}),$ (2)
where $\hat{\mathbf{I}}\in\mathbb{R}^{\hat{D}\times\hat{H}\times\hat{W}\times\hat{d}_{\text{I}}}$ is the embedded 3D features for surrounding camera images with the similar dimensions as $\hat{\mathbf{P}}$ except for $\hat{d}_{\text{I}}$ feature channels. With $\hat{\mathbf{P}}$ and $\hat{\mathbf{I}}$, we further concatenate them along feature dimension and apply the fusion encoder $f^{\text{enc}}_{\text{fusion}}$ to get the fusion feature $\hat{\mathbf{F}}\in\mathbb{R}^{\hat{D}\times\hat{H}\times\hat{W}\times\hat{d}_{\text{F}}}$,
@@
Loss Function. To guide $f^{\text{enc}}_{\text{P}}$, $f^{\text{enc}}_{\text{I}}$ and $f^{\text{enc}}_{\text{fusion}}$ to learn good representations in an unsupervised manner, CLAP first embed the fusion features $\hat{\mathbf{F}}$ with a shallow 3D convolution network $f^{\text{3D}}$ to get $\tilde{\mathbf{F}}=f^{\text{3D}}(\hat{\mathbf{F}})$ and we have $\tilde{\mathbf{F}}\in\mathbb{R}^{\hat{D}\times\hat{H}\times\hat{W}\times\hat{d}_{\text{F}}}$. Then a rendering loss $\mathcal{L}_{\text{rend}}$ for masked-reconstruction is applied on $\tilde{\mathbf{F}}$. Furthermore, a swapping prototype assignment prediction scheme $\mathcal{L}_{\text{proto}}$ is utilized in order to incorporate interaction of the two modalities into pre-training. The overall loss function is as below:
$\mathcal{L}=\mathcal{L}_{\text{rend}}(\mathbf{P},\tilde{\mathbf{F}},\mathcal{T})+\omega_{\text{proto}}\times\mathcal{L}_{\text{proto}}(\hat{\mathbf{P}},\hat{\mathbf{I}}),$ (4)
with $\omega_{\text{proto}}$ as a weighting parameter to balance the losses.

### v2

@@ Abstract
Unsupervised 3D representation learning reduces the burden of labeling multimodal 3D data for fusion perception tasks. Among different pre-training paradigms, differentiable-rendering-based methods have shown most promise. However, existing works separately conduct pre-training for each modalities due to computational costs of processing large point clouds with images. As such, mutual benefit of high-level semantics (from image) and 3D structure (from point cloud) has not been exploited. To address this gap, we propose a joint unsupervised differentiable-rendering-based pre-training method for images and point clouds, termed CLAP, short for Curvature sampLing and leArnable Prototype. Specifically, our method overcomes the computational hurdle by Curvature Sampling to select the more informative points/pixels for pre-training. To uncover the performance benefits brought by their complementarity, we propose to use learnable prototypes to represent parts of the 3D scenes in a common feature space and an Expectation-Maximization training scheme to associate embeddings of each modality to prototypes. We further propose a swapping prediction loss that explores their interplay through prototypes along with a Gram Matrix Regularization term to maintain training stability. Experiments on NuScenes and Waymo datasets show that CLAP achieves up to $100\%$ more performance gain as compared to previous SOTA pre-training methods. Codes and models will be released.
@@ 1 Introduction
3D perception facilitates spatial applications such as autonomous driving. The autonomous system, on which these applications are deployed, are typically equipped with multiple sensors including visual ones like cameras that produce RGB images, and range sensors like LiDAR (Light-Detection-And-Ranging) that generate point clouds. Fusion of these two modalities [30, 26, 22, 25, 21, 32, 8] have generally improved over use of a single modality, e.g., camera [11, 7, 43, 50, 60, 20] or LiDAR [39, 55, 42, 59, 12, 1] separately, in 3D perception performance.
@@
Figure 1: Unlike previous SOTA unsupervised 3D representation learning method UniPAD [58] that separately pre-train LiDAR and camera encoders with differential rendering (a), our proposed method CLAP conducts joint pre-training for fusion perception.
However, training multimodal 3D perception models is expensive as labeling in 3D space is notoriously time-and-energy-consuming. Unsupervised 3D representation learning, which pre-trains backbones without any label and fine-tuned the pre-trained weight for downstream performance improvement, has shown the potential to alleviate the labeling burden in 3D perception. Amongst the many unsupervised 3D pre-training methods [24, 19, 6, 57, 53, 18, 62, 58], use of mask auto-encoding (reconstruction) and differntiable rendering has emerged as the most performant. However, encoding and processing high-dimensional multi-modal data (images and point clouds) is computationally expensive: If one were to pre-train using all points and pixels within the point cloud and image, respectively, even the most advanced GPU to date (H100) is only able to hold a batch size of 1. Therefore, existing methods, e.g., UniPad [58], has conventionally pre-trained each modality separately. While the two modalities are complementary (semantics from image, and structure from point clouds), current methods are unable to exploit their mutual benefits during pre-training.
To address this pain point, we propose a joint unsupervised pre-training method of image and point cloud modalities based on differentiable rendering. Our method, CLAP, short for Curvature sampLing and leArnable Prototype, addresses the primary computational challenge by a curvature sampling strategy, which comes from the observation that there exists redundancy in information when sampling multiple points on the same flat surface. This is enabled by estimating the curvature of each point in the 3D space by taking second order derivative of the SDF (signed distance field) function. This allows us to identify points at which large change in the surface occurs. Through calibration, we can obtain the corresponding pixel within the image, around which we sample assuming the coordinates are normally distributed. This sampling strategy captures variations in the point cloud and in turn provide more informative points than the conventional sampling strategy in [58].
As our curvature sampling strategy has reduced the computational burden to allow both modalities to be simultaneously processed, we propose to model the interplay between image and point cloud modalities through a set of learnable prototypes, which represents parts of the 3D scene and enables a common feature space to bridge the two modalities. These prototypes are trained via an Expectation-Maximization (EM) algorithm that maximizes similarity between embeddings for each modality and the set of prototypes. Furthermore, we propose to use the swapping prediction loss to explore the interaction between the two modalities. Last but not least, we utilize a Gram Matrix Regularization term to minimize similarity across prototypes, which addresses collapse of prototype when trained naively.
Our contributions are 4-folds: (1) We propose a curvature sampling strategy to identify informative points (and pixels) for sampling, which enables the first joint differentiable-rendering-based pre-training method for fusion perception. (2) Learnable prototypes are utilized to learn a common feature space and an Expectation-Maximization approach is proposed to train the prototypes to represent parts of the 3D scene. (3) We further propose to use a swapping prediction loss for modality interplay and a Gram Matrix Regularization loss that avoids collapse of prototype learning. (4) Through extensive experiments on the popular autonomous driving datasets NuScenes [3] and Waymo [45], we demonstrate the effectiveness of CLAP . For example, CLAP achieves up to $100\%$ more improvement than previous SOTA 3D pre-training methods and shows potential scaling property.
@@ 2 Related Work
Fusion 3D Object Detection. Light-Detection-And-Ranging (LiDAR) and camera are important sensors for autonomous driving perception. Previous works mainly focus on single-modality 3D perception. For LiDAR-based 3D object detection, there are three main streams with different embedding schemes for point clouds inputs. 1) Point-based methods [39, 41] utilize point-level embeddings for 3D object detection. 2) Voxel-based methods [55, 10, 59, 1, 56, 12] voxelize the 3D scene and use sparse convolution or transformer for embedding. 3) Point-voxel-combined methods [40, 42] utilize both embeddings from 1) and 2). For camera-based 3D perception, [11, 7, 43, 50, 60, 20] embed 2D features on image plane and project these 2D features into 3D space with estimated depth. Recently, research starts to focus on fusion 3D perception [8, 30, 26, 22, 25, 21, 32] with camera images and LiDAR point clouds as inputs. These methods mainly focus on integrating embeddings of different modalities and train both the 2D and 3D backbones in an supervised manner. As labeling in 3D space is costly, we explore unsupervised 3D representation learning for fusion 3D perception.
@@
3D Pre-training. Annotating for 3D data is notoriously time- and energy-consuming and the emergence of unsupervised representation learning for 2D image [15, 47, 4, 13, 51, 16] provides a promising way to alleviate the annotation burden. Existing works [52, 27, 17, 6, 19, 24, 28, 38, 36, 57, 53, 18, 62] in unsupervised 3D representation learning into scene-level 3D point clouds can be divided into two contrastive-based and masked-and-reconstruction-based paradigms. Contrastive-based works [52, 27, 17, 6, 19, 24, 36, 28] propose various ways to build suitable views and conduct contrastive learning to improve the performance in downstream perception task. Inspired by [16] in image domain, [57, 53, 18, 62] propose to first mask the input point clouds and pre-train the 3D encoders with a shallow decoder for reconstructing the unmasked inputs. [44, 61, 23] are pioneering works to introduce contrastive learning into fusion perception. They consider camera and LiDAR embeddings as different views of the scene and apply the contrastive loss between the two modalities. Inspired by the success of neural field in representing 3D scenes [33, 49] and previous attempts to introduce neural rendering to 3D pre-training for point clouds [54, 18, 62], UniPAD [58] proposes to use a differentiable-rendering decoder for masked-and-reconstruction pre-training and achieves SOTA performance for unsupervised 3D representation learning on fusion 3D perception. However, due to the high GPU memory consumption, UniPAD [58] is only able to separately pre-train the image and point cloud encoders and fails to utilize the interaction between modalities during pre-training. In this paper, we explore unsupervised joint pre-training for 2D and 3D backbones via differentiable rendering with Curvature Sampling and Prototype Learning. Previous work like SwAV [4] is related to our prototype learning part. SwAV [4] explores learnable prototypes to represent different categories in the same modality (image) and uses swapping assignment prediction loss for different views of the same image instance. On the contrary, learnable prototypes in CLAP are used to represent part of the 3D scenes and learn the interaction between modalities, which differs from the context in [4]. Thus we propose a Expectation-Maximization training scheme to maximize similarity between prototypes and 3D embeddings. Then a swapping prediction loss is used to learn the modality interaction. Furthermore, to avoid prototypes collapsing to the same vector, we propose a Gram Matrix Regularization loss.
@@ 3 Method
Figure 2: The pipeline of CLAP. In order to jointly pre-train the LiDAR, camera and fusion encoders, we first embed the paired LiDAR point clouds and camera images with $f^{\text{enc}}_{\text{P}}$, $f^{\text{enc}}_{\text{I}}$ and $f^{\text{enc}}_{\text{fusion}}$. Then based on the fusion features, CLAP applies differentiable rendering to predict both depth and rgb with the SDF and RGB values of the sampled points along LiDAR/camera rays from $f^{\text{SDF}}$ and $f^{\text{RGB}}$, with which we compute loss against the observed LiDAR point cloud and camera images. To make joint pre-training feasible, we propose Curvature Sampling to sample informative parts of the 3D scene, as descriped in (a). Furthermore, we propose to use learnable prototypes to represent parts of objects in a common feature space and utilize an Expectation-Maximization approach to maximize the similarity between prototypes and 3D embeddings of each modality. To delve deeper into the interplay of image semantics and LiDAR geometry, we use swapping prototype prediction loss. Finally, we propose a Gram Matrix Regularization loss to prevent collapse of prototype learning.
In this section, we introduce CLAP for joint unsupervised 3D pre-training via differentiable rendering on fusion 3D perception. As described in Fig. 2, CLAP pre-trains the image, LiDAR and fusion encoders jointly with Neural Field Rendering. In order to enable joint pre-training, Curvature Sampling is proposed as shown in (a) of Fig. 2. To further make use of both modalities, we utilize learnable prototypes to represent parts of the 3D scenes as shown in Fig. 2 (b). To optimize the learnable prototype, we train a common feature space with an Expectation-Maximization training scheme and incorporate interaction between modalities by a swapping prediction loss. Finally a Gram Matrix Regularization loss is proposed to avoid collapse in prototype learning. We first discuss the formulation and overall pipeline in Section 3.1. Then we introduce the details about neural field and differentiable rendering in Section 3.2. Finally, we describe the curvature sampling and prototype learning separately in Section 3.3 and 3.4.
@@ 3.1 Formulation and Pipeline
Notations. To begin with, we denote the input image set from $N_{\text{cam}}$ cameras as =$\mathcal{I}=\{\mathbf{I}_{n}\in\mathbb{R}^{H\times W\times 3}\}_{n=1}^{N_{\text{cam}}}$ and LiDAR point cloud as $\mathbf{P}\in\mathbb{R}^{N_{p}\times(3+d)}$. $H$ and $W$ are the height and width of the images and each pixel on the images has $3$ values for RGB. $N_{p}$ is the number of points in the LiDAR point cloud and each of them contains $xyz$-location and $d$ feature channels. For example, in NuScenes [3] dataset, $d=2$ represents the intensity and timestamp of each point and there are $N_{\text{cam}}=6$ surrounding cameras on the autonomous vehicle. For each pair of camera image and LiDAR point cloud, we have the transformation matrix $\mathbf{T}_{n}\in\mathbb{R}^{3\times 4}$ indicating the projection between the camera plane and LiDAR coordinate, where $n=1,2,...,N_{\text{cam}}$.
@@
where $\hat{\mathbf{P}}\in\mathbb{R}^{\hat{D}\times\hat{H}\times\hat{W}\times\hat{d}_{\text{P}}}$ is the embedded 3D features for LiDAR point cloud. $\hat{D}$, $\hat{H}$ and $\hat{W}$ are spatial resolutions of the embedded features and $\hat{d}_{\text{P}}$ is number of feature channels after encoding. Then for camera images $\mathcal{I}$, $f^{\text{enc}}_{\text{I}}$ encodes them with swin transformer [29] and uses $\mathcal{T}=\{\mathbf{T}_{n}\}_{n=1}^{N_{\text{cam}}}$ to project the 2D features to 3D space.
$\hat{\mathbf{I}}=f^{\text{enc}}_{\text{I}}(\mathcal{I},\mathcal{T},\mathbf{P}),$ (2)
where $\hat{\mathbf{I}}\in\mathbb{R}^{\hat{D}\times\hat{H}\times\hat{W}\times\hat{d}_{\text{I}}}$ is the embedded 3D features for surrounding camera images with the similar dimensions as $\hat{\mathbf{P}}$ except for $\hat{d}_{\text{I}}$ feature channels. The projection of 2D features to 3D space is similar to [30]: we transform LiDAR points back to image planes with $\mathcal{T}$ and use the projected depths to project the 2D features to 3D space. With $\hat{\mathbf{P}}$ and $\hat{\mathbf{I}}$, we further concatenate them along feature dimension and apply the fusion encoder $f^{\text{enc}}_{\text{fusion}}$ to get the fusion feature $\hat{\mathbf{F}}\in\mathbb{R}^{\hat{D}\times\hat{H}\times\hat{W}\times\hat{d}_{\text{F}}}$ with $\hat{d}_{\text{F}}$ feature dimensions,
@@
Loss Function. To guide $f^{\text{enc}}_{\text{P}}$, $f^{\text{enc}}_{\text{I}}$ and $f^{\text{enc}}_{\text{fusion}}$ to learn good representations in an unsupervised manner, CLAP first embed the fusion features $\hat{\mathbf{F}}$ with a shallow 3D convolution network $f^{\text{3D}}$ to get $\tilde{\mathbf{F}}=f^{\text{3D}}(\hat{\mathbf{F}})$ and we have $\tilde{\mathbf{F}}\in\mathbb{R}^{\hat{D}\times\hat{H}\times\hat{W}\times\hat{d}_{\text{F}}}$. Then a rendering loss $\mathcal{L}_{\text{rend}}$ is applied on $\tilde{\mathbf{F}}$ for masked-reconstruction on both point clouds $\mathbf{P}$ and images $\mathcal{I}$. Furthermore, a prototype learning scheme $\mathcal{L}_{\text{proto}}$ is utilized in order to bridge the two modalities and incorporate interaction of image semantics and LiDAR geometry into pre-training. The overall loss function is as below:
$\mathcal{L}=\omega_{\text{r}}\times\mathcal{L}_{\text{rend}}(\mathbf{P},\tilde{\mathbf{F}},\mathcal{I})+\omega_{\text{proto}}\times\mathcal{L}_{\text{proto}}(\hat{\mathbf{P}},\hat{\mathbf{I}}),$ (4)
with $\omega_{\text{r}}$ and $\omega_{\text{proto}}$ as weighting parameters to balance the losses.